In triangle ABC, angle B is 4 times A and angle C is 12 degrees less than 5 times angle A. Find the size of the angles.
The size of angle A is ___ degrees?
The size of angle B is ___ degrees?
The size of angle C is ___ degrees?
Simplify your answer for each. Type an integer or a decimal.

Found 2 solutions by swincher4391, stanbon:
Answer by swincher4391(1107)

We know that A + B +C = 180
They tell us:
B = 4A
C = 5A - 12
Then A + 4A + (5A-12) = 180
10A - 12 = 180
10A = 192
A = 19.2º
B = 4(19.2) = 76.8º
19.2 + 76.8 + C = 180
96 + C = 180
C = 180 - 96
C = 84º

The angles are 19.2, 76.8, 84.
What you asked for, Angle A, is 19.2º.
Hope this helped!

Answer by stanbon(75887)

In triangle ABC, angle B is 4 times A and angle C is 12 degrees less than 5 times angle A. Find the size of the angles.
-------------
Equations:
A + B + C = 180
B = 4A
C = 5A-12
------
Substitute and solve for "A":
A + 4A + 5A-12 = 180
10A = 192
A = 19.2 degrees
B = 4A = 76.8 degrees
C = 5A-12 = 84 degrees
==========================
Cheers,
Stan H.
==================



The size of angle A is ___ degrees?
The size of angle B is ___ degrees?
The size of angle C is ___ degrees?